OpenStudy (anonymous):

Is f(x) = sin^2(x) an even, odd, or neither function. Why? I got even due to symmetry, but I am not sure.

OpenStudy (anonymous):

sin(x) sin(x) change it to negative sin(-x) sin(-x)= -sin(x) - sin(x)= sin^2(x) even

OpenStudy (anonymous):

f is even if f(-x)=f(x) f is odd if f(-x)=-f(x)